# Continuous Input Embedding Size Search For Recommender Systems

### v5

@@ 1. Introduction
In light of these challenges, we propose continuous input embedding size search (CIESS), which is an RL-based algorithm that can efficiently operate in a (near [^1] This is due to the fact that embedding sizes can only be integers.) continuous action space for embedding size search. To enable generalization to a continuous action space, we build CIESS upon an actor-critic paradigm (Grondman et al., 2012), specifically the twin delayed deep deterministic policy gradient (TD3) (Fujimoto et al., 2018), where we have designed a policy/actor network that can determine the best embedding size from an interval based on the state of each user/item. Compared with existing embedding size search counterparts (Joglekar et al., 2020; Zhao et al., 2021b; Liu et al., 2020) that only support discrete embedding size search from a small pool of actions, this is the first work to explore a large, continuous, and fine-grained RL search space of embedding sizes, thus unlocking the full potential of learning a compact recommender system with optimal recommendation performance. Furthermore, CIESS is a versatile embedding size search approach that does not hold any assumptions on the backbone recommendation model, and is compatible with a variety of latent factor recommenders that require an embedding table.
@@ 2.1. Base Recommender with Masked Embeddings
With the mask $\mathbf{M}$, for each user/item, we can retain the first $d_{n}$ elements of its full embedding while setting all succeeding dimensions to 0. It is worth noting that, performing embedding sparsification by masking unwanted dimensions with zeros is a commonly adopted approach in lightweight recommender systems (Liu et al., 2021; Lyu et al., 2022), as the resulting embedding table can take advantage of the latest sparse matrix storage techniques (Virtanen et al., 2020; Sedaghati et al., 2015) that bring negligible cost for storing zero-valued entries.
@@ 2.2.2. State
The state $s$ is the input to the policy network (i.e., actor in CIESS) that drives the decision-making on each user/item-specific embedding size. (Zhao et al., 2021b) shows that the popularity (i.e., the number of interactions) and the current embedding size decision of the user/item are effective in providing the policy network the context for subsequent search. Our method inherits this design, with an additional quality indicator $a$ that records the recommendation accuracy fluctuation under the current policy:
@@
where $f_{n}$ is the popularity of the user/item, normalized by the corresponding maximum/minimum frequency among all users $f_{max}=\max_{u\in\mathcal{U}}(f_{u})$$f_{min}=\min_{u\in\mathcal{U}}(f_{u})$ or items $f_{max}=\max_{v\in\mathcal{V}}(f_{v})$$f_{min}=\min_{v\in\mathcal{V}}(f_{v})$, which are observed from training data. $d_{n}$ is the current embedding size allocated to a user/item, and $q_{n}$ quantifies the changes in the recommendation quality when the embedding size decreases from $d_{max}$ to the current $d_{n}$. Compared with (Zhao et al., 2021b), incorporating this quality indicator into the state is able to help trace the impact from the most recent action (i.e., embedding size) to the recommendation effectiveness, which can encourage the policy network to better balance the memory cost and performance with the embedding sizes selected from a vast, continuous action space.
@@ 2.2.5. Actor and Critic
Action Exploration with Random Walk. As discussed in Section 1, such value-based RL optimization heavily relies on learning an accurate critic, which can hardly be guaranteed in continuous embedding size search tasks with numerous actions and states. Although TD3 additionally trains an independent critic network (two additional user/item critic networks in our case) and uses the minimum of their outputs as the predicted Q-value to lower the estimation variance, such a conservative strategy may result in underestimation of the Q-value and eventually suboptimal performance (Liu et al., 2023). Therefore, a common practice is to adopt exploration strategies on possible actions. A typical example is the greedy strategy that always selects the action with highest Q-value (Dulac-Arnold et al., 2015), which can be extremely inefficient to obtain in our large action space. Hence, in addition to the exploration noise in Eq.(11), we further propose a sample-efficient action exploration strategy based on random walk, as shown in Figure 2. To avoid evaluating all possible actions at each state as in the greedy strategy, we only compute Q-values on a sequence of actions mutated from the computed $\hat{d}^{i}_{n}$, where the action returning the highest Q-value will be selected. The rationale is, for a given user/item, if we keep fine-tuning its embedding size based on currently the best one $\hat{d}^{i}_{n}$ predicted by the actor, there will be a higher chance that we can obtain a more performant embedding size choice. Specifically, to efficiently identify the sequence of actions for exploration, we construct a graph of actions where each action $d$ is connected to a set of neighbor actions $\mathcal{A}_{d}$. Since each action is an integer-valued embedding size, we can easily measure the distance between any two actions by taking the absolute value of their arithmetic difference. For each action $d$, the distance to any of its neighbors are below a predefined threshold $t$:
@@ 2.3. Selective Retraining for Sparsified Embeddings
We put together a pseudo code for CIESS in Algorithm 1. The embedding size search policy is trained with RL for $M$ episodes, where each episode iterates for $T$ times. In each iteration, CIESS performs random walk-based action exploration and decides the embedding sizes for all users and items (lines 6-9), train the recommender $F_{\Theta}(\cdot)$ with the embedding sizes selected to obtain the instant reward $r^{i}_{n}$ and next iteration’s state $s^{i+1}_{n}$ (lines 10-11). The transition described in Section 2.2.4 is appended to the replay buffer $\mathcal{B}$ to facilitate optimization of the critic and actor networks in the search function (lines 12-14). We omit the twin network-based optimization process in TD3 and set a pointer to the original paper (Fujimoto et al., 2018) for finer details. Notably, as a fresh recommender $F_{\Theta}(\cdot)$ needs to be trained from scratch for every iteration’s embedding size decision, we restrict the training process (line 10) to only run for a specified number of epochs (5 in our case) over $\mathcal{D}$ to ensure efficiency. By fixing all hyperparameters and only varying $\mathbf{M}$ for $F_{\Theta}(\cdot)$, this offers us sufficient confidence in comparing different embedding sizes’ performance potential without an excessively time-consuming process. We hereby introduce how to obtain a fully trained recommender under a specific memory budget, which corresponds to line 17 in Algorithm 1.
@@ 3.1. Base Recommenders and Comparative Methods
  • OptEmb (Lyu et al., 2022): It trains a supernet to learn field-wise thresholds, and then uses evolutionary search to derive the optimal embedding size for each field.
@@ 3.2. Evaluation Protocols
For effectiveness, we adopt Recall@$k$ and NDCG@$k$ as our metrics by setting $k\in\{5,20\}$. For CIESS, PEP, ES, and MR, we test the recommendation performance under three sparsity ratios $c\in\{80\%,90\%,95\%\}$. For each of these four methods, it is guaranteed that the compressed embedding table has no more than $cd_{max}\times(|\mathcal{U}|+|\mathcal{V}|)$ usable parameters, where the full embedding size $d_{max}=128$. Notably, since ESAPN and OptEmb have a more performance-oriented design and do not offer a mechanism to precisely control the resulted embedding sparsity, we only report the performance achieved by their final embedding tables.
@@ 3.4. Comparison of Overall Performance
Meanwhile, ESAPN and OptEmb have respectively resulted in a 75% and 79% sparsity rate on average, and failed to meet the lowest 80% target most of the time. Although both have retained relatively more parameters than all other methods in many test cases, their recommendation performance constantly falls behind CIESS. For example, when paired with LightGCN on MovieLens-1M dataset, ESAPN needs to consume more parameters (72% embedding sparsity) to obtain a competitive Recall@5 score w.r.t. CIESS under 80% sparsity. It is also worth noting that, on Yelp2018 dataset, CIESS under 90% sparsity even outperforms ESAPN and OptEmb with a much lower sparsity (e.g., 67% for OptEmb 76% for ESAPN). In short, with the same memory consumption, CIESS delivers stronger performance; and at the same performance level, CIESS is more memory-efficient. Hence, the results showcase the continuous embedding size search in CIESS is more advantageous in preserving recommendation accuracy.
@@ 4.1. Latent Factor Recommenders
Designing deep recommender models heavily relies on the expertise of professionals. To alleviate the need for human engagement, Automated Machine Learning for Recommender Systems has been studied to automate the design process for recommender models in a task-specific and data-driven manner. So far several research directions including feature selection search (Wang et al., 2022; Luo et al., 2019), embedding dimension search (Liu et al., 2021, 2020; Zhao et al., 2021b; Qu et al., 2022), feature interaction search (Su et al., 2021, 2022), model architecture search (Song et al., 2020; Cheng et al., 2022), and other components search (Zhao et al., 2021a; Meng et al., 2021) have been proposed (Zheng et al., 2022). The first kind reduces computation cost by filtering feature fields based on learnable importance scores (Wang et al., 2022; Luo et al., 2019). The second line of works proposes dynamic embedding sizes for each feature (Liu et al., 2020, 2021; Zhao et al., 2021b; Qu et al., 2022). The third kind prevents recommender models from enumerating all high-order feature interactions when learning the interactions between features (Su et al., 2021, 2022). Model architecture search models explores different network architectures and determines the optimal architecture (Song et al., 2020; Cheng et al., 2022). Other components search focuses on optimize the loss function and feature interaction function (Zhao et al., 2021a; Meng et al., 2021).
@@ 5. Conclusion
Latent factor recommenders use vectorized embeddings with a single and uniform size, leading to substandard performance and excessive memory complexity. To overcome this issue, we proposed an RL-based, model-agnostic embedding size search algorithm, CIESS, that can select tailored embedding size from a continuous interval for each user/item, thus refining the representation expressiveness with minimal memory costs. In the future, we will investigate embedding size search for streaming services, which allows for more usable information for constructing the RL states.

### v6

@@ 1. Introduction
In light of these challenges, we propose continuous input embedding size search (CIESS), which is an RL-based algorithm that can efficiently operate in a (near [^1] This is due to the fact that embedding sizes can only be integers.) continuous action space for embedding size search. To enable generalization to a continuous action space, we build CIESS upon an actor-critic paradigm (Grondman et al., 2012), specifically the twin delayed deep deterministic policy gradient (TD3) (Fujimoto et al., 2018), where we have designed a policy/actor network that can determine the best embedding size from an interval based on the state of each user/item. Compared with existing embedding size search counterparts (Joglekar et al., 2020; Zhao et al., 2021a; Liu et al., 2020) that only support discrete embedding size search from a small pool of actions, this is the first work to explore a large, continuous, and fine-grained RL search space of embedding sizes, thus unlocking the full potential of learning a compact recommender system with optimal recommendation performance. Furthermore, CIESS is a versatile embedding size search approach that does not hold any assumptions on the backbone recommendation model, and is compatible with a variety of latent factor recommenders that require an embedding table.
@@ 2.1. Base Recommender with Masked Embeddings
With the mask $\mathbf{M}$, for each user/item, we can retain the first $d_{n}$ elements of its full embedding while setting all succeeding dimensions to 0. It is worth noting that, performing embedding sparsification by masking unwanted dimensions with zeros is a commonly adopted approach in lightweight recommender systems (Liu et al., 2021; Lyu et al., 2022), as the resulting embedding table can take advantage of the latest sparse matrix storage techniques (Virtanen and others, 2020; Sedaghati et al., 2015) that bring negligible cost for storing zero-valued entries.
@@ 2.2.2. State
The state $s$ is the input to the policy network (i.e., actor in CIESS) that drives the decision-making on each user/item-specific embedding size. (Zhao et al., 2021a) shows that the popularity (i.e., the number of interactions) and the current embedding size decision of the user/item are effective in providing the policy network the context for subsequent search. Our method inherits this design, with an additional quality indicator $a$ that records the recommendation accuracy fluctuation under the current policy:
@@
where $f_{n}$ is the popularity of the user/item, normalized by the corresponding maximum/minimum frequency among all users $f_{max}=\max_{u\in\mathcal{U}}(f_{u})$$f_{min}=\min_{u\in\mathcal{U}}(f_{u})$ or items $f_{max}=\max_{v\in\mathcal{V}}(f_{v})$$f_{min}=\min_{v\in\mathcal{V}}(f_{v})$, which are observed from training data. $d_{n}$ is the current embedding size allocated to a user/item, and $q_{n}$ quantifies the changes in the recommendation quality when the embedding size decreases from $d_{max}$ to the current $d_{n}$. Compared with (Zhao et al., 2021a), incorporating this quality indicator into the state is able to help trace the impact from the most recent action (i.e., embedding size) to the recommendation effectiveness, which can encourage the policy network to better balance the memory cost and performance with the embedding sizes selected from a vast, continuous action space.
@@ 2.2.5. Actor and Critic
Action Exploration with Random Walk. As discussed in Section 1, such value-based RL optimization heavily relies on learning an accurate critic, which can hardly be guaranteed in continuous embedding size search tasks with numerous actions and states. Although TD3 additionally trains an independent critic network (two additional user/item critic networks in our case) and uses the minimum of their outputs as the predicted Q-value to lower the estimation variance, such a conservative strategy may result in underestimation of the Q-value and eventually suboptimal performance (Liu et al., 2023). Therefore, a common practice is to adopt exploration strategies on possible actions. A typical example is the greedy strategy that always selects the action with highest Q-value (Dulac-Arnold et al., 2015), which can be extremely inefficient to obtain in our large action space. Hence, in addition to the exploration noise in Eq.(11), we further propose a sample-efficient action exploration strategy based on random walk, as shown in Figure 2. To avoid evaluating all possible actions at each state as in the greedy strategy, we only compute Q-values on a sequence of actions mutated from the computed $\hat{d}^{i}_{n}$, where the action returning the highest Q-value will be selected. The rationale is, for a given user/item, if we keep fine-tuning its embedding size based on currently the best one $\hat{d}^{i}_{n}$ predicted by the actor, there will be a higher chance that we can obtain a more performant embedding size choice. Specifically, to efficiently identify the sequence of actions for exploration, we construct a graph of actions where each action $d$ is connected to a set of neighbor actions $\mathcal{A}_{d}$. Since each action is an integer-valued embedding size, we can easily measure the distance between any two actions by taking the absolute value of their arithmetic difference. For each action $d$, the distance to any of its neighbors is below a predefined threshold $t$:
@@ 2.3. Selective Retraining for Sparsified Embeddings
We put together a pseudo code for CIESS in Algorithm 1. The embedding size search policy is trained with RL for $M$ episodes, where each episode iterates for $T$ times. In each iteration, CIESS performs random walk-based action exploration and decides the embedding sizes for all users and items (lines 6-9), trains the recommender $F_{\Theta}(\cdot)$ with the embedding sizes selected to obtain the instant reward $r^{i}_{n}$ and next iteration’s state $s^{i+1}_{n}$ (lines 10-11). The transition described in Section 2.2.4 is appended to the replay buffer $\mathcal{B}$ to facilitate optimization of the critic and actor networks in the search function (lines 12-14). We omit the twin network-based optimization process in TD3 and set a pointer to the original paper (Fujimoto et al., 2018) for finer details. Notably, as a fresh recommender $F_{\Theta}(\cdot)$ needs to be trained from scratch for every iteration’s embedding size decision, we restrict the training process (line 10) to only run for a specified number of epochs (5 in our case) over $\mathcal{D}$ to ensure efficiency. By fixing all hyperparameters and only varying $\mathbf{M}$ for $F_{\Theta}(\cdot)$, this offers us sufficient confidence in comparing different embedding sizes’ performance potential without an excessively time-consuming process. We hereby introduce how to obtain a fully trained recommender under a specific memory budget, which corresponds to line 17 in Algorithm 1.
@@ 3.1. Base Recommenders and Comparative Methods
  • OptEmbed (Lyu et al., 2022): It trains a supernet to learn field-wise thresholds, and then uses evolutionary search to derive the optimal embedding size for each field.
@@ 3.2. Evaluation Protocols
For effectiveness, we adopt Recall@$k$ and NDCG@$k$ as our metrics by setting $k\in\{5,20\}$. For CIESS, PEP, ES, and MR, we test the recommendation performance under three sparsity ratios $c\in\{80\%,90\%,95\%\}$. For each of these four methods, it is guaranteed that the compressed embedding table has no more than $cd_{max}\times(|\mathcal{U}|+|\mathcal{V}|)$ usable parameters, where the full embedding size $d_{max}=128$. Notably, since ESAPN and OptEmbed have a more performance-oriented design and do not offer a mechanism to precisely control the resulted embedding sparsity, we only report the performance achieved by their final embedding tables.
@@ 3.4. Comparison of Overall Performance
Meanwhile, ESAPN and OptEmbed have respectively resulted in a 75% and 79% sparsity rate on average, and failed to meet the lowest 80% target most of the time. Although both have retained relatively more parameters than all other methods in many test cases, their recommendation performance constantly falls behind CIESS. For example, when paired with LightGCN on MovieLens-1M dataset, ESAPN needs to consume more parameters (72% embedding sparsity) to obtain a competitive Recall@5 score w.r.t. CIESS under 80% sparsity. It is also worth noting that, on Yelp2018 dataset, CIESS under 90% sparsity even outperforms ESAPN and OptEmbed with a much lower sparsity (e.g., 67% for OptEmbed 76% for ESAPN). In short, with the same memory consumption, CIESS delivers stronger performance; and at the same performance level, CIESS is more memory-efficient. Hence, the results showcase the continuous embedding size search in CIESS is more advantageous in preserving recommendation accuracy.
@@ 4.1. Latent Factor Recommenders
Designing deep recommender models heavily relies on the expertise of professionals. To alleviate the need for human engagement, Automated Machine Learning for Recommender Systems has been studied to automate the design process for recommender models in a task-specific and data-driven manner. So far several research directions including feature selection search (Wang et al., 2022; Luo et al., 2019), embedding dimension search (Liu et al., 2021, 2020; Zhao et al., 2021a; Qu et al., 2022), feature interaction search (Su et al., 2021, 2022), model architecture search (Song et al., 2020; Cheng et al., 2022), and other components search (Zhao et al., 2021b; Meng et al., 2021) have been proposed (Zheng et al., 2022). The first kind reduces computation cost by filtering feature fields based on learnable importance scores (Wang et al., 2022; Luo et al., 2019). The second line of works proposes dynamic embedding sizes for each feature (Liu et al., 2020, 2021; Zhao et al., 2021a; Qu et al., 2022). The third kind prevents recommender models from enumerating all high-order feature interactions when learning the interactions between features (Su et al., 2021, 2022). Model architecture search models explores different network architectures and determines the optimal architecture (Song et al., 2020; Cheng et al., 2022). Other components search focuses on optimize the loss function and feature interaction function (Zhao et al., 2021b; Meng et al., 2021).
@@ 5. Conclusion
Latent factor recommenders use vectorized embeddings with a single and uniform size, leading to substandard performance and excessive memory complexity. To overcome this issue, we proposed an RL-based, model-agnostic embedding size search algorithm, CIESS, that can select tailored embedding size from a continuous interval for each user/item, thus refining the representation expressiveness with minimal memory costs. Future work could explore richer and more robust importance modeling by incorporating signals such as weight magnitudes (Janowsky, 1989; Ström, 1997; Thimm and Fiesler, 1995) or model confidence (Qu et al., 2021, 2021). Such extensions could improve capacity allocation for rare-but-informative entities and mitigate popularity bias.